OpenStudy (anonymous):

trig equation tan ()=-1

OpenStudy (adamaero):

This will most likely be in the front or back cover of your textbook.

OpenStudy (adamaero):

If you haven't memorized it: arctan(1) = pi/4

OpenStudy (adamaero):

And since the negative goes right through for tan(#), and sin(#) too attach a negative to that pi/4

OpenStudy (adamaero):

Funny, you posted. I was doing calculus homework and needed the arctan of this for the bounds of an integral.

OpenStudy (anonymous):

Tan=sin/cos pi/4=45degree (rad2/2/rad2/2)=1 -1 cause arch